# Estimation of AMOC transition probabilities using a machine learning based rare-event algorithm

### v1

@@ 3 Results: F-transitions
In Castellana et al. (2019), the score function to determine the probability of occurrence of a F-transition of the AMOC was prescribed as:
@@
where $\mathbf{x}\in\mathbb{R}^{d}$ is a state in the phase space and the exponent $ON$ refers to the steady on-state of the AMOC model. $\Phi^{F}_{S}(\mathbf{x})$ is equal to $1$ whenever $q_{N}(\mathbf{x})=0$. Here, we define the set $A$ to be a neighborhood of $q^{ON}_{n}$ such that $q_{N}(\mathbf{x})\geq q_{S}(\mathbf{x})$. Given a certain parameter $\rho$, we define $A=\{\mathbf{x}\in\mathbb{R}^{n}\ |\ ||\mathbf{x}-\mathbf{x}^{ON}||\leq\rho\ \&\ q_{N}(\mathbf{x})\geq q_{S}(\mathbf{x})\}$. $\Phi^{F}_{S}(\mathbf{x})$ is set to 00 in this area.
The score function estimated by the RC is called $\Phi_{R}^{F}$. In what follows, the combination of TAMS with the prescribed score function $\Phi^{F}_{S}$ is called TAMS-$S$, and the combination of TAMS with $\Phi_{R}^{F}$ is called TAMS-$R$. We consider $100$ equally spaced values of $\overline{E_{a}}\in[0.06,0.35]$ and $f_{\sigma}\in[0.005,0.5]$. TAMS uses an ensemble of $N=1,000$ trajectories, $n_{c}=10$ of which are discarded at each iteration; $T_{\mathrm{max}}=100$ years. For each couple $(\overline{E_{a}},f_{\sigma})$, TAMS-$S$ and TAMS-$R$ are run both $M=30$ times.
@@ 3.1 Consistency indicator $\mathcal{C}$
First, we must make sure that the transition probability estimates obtained with both methods are consistent. We call $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ the mean probability estimates, respectively, obtained with TAMS-$S$ and TAMS-$R$. We call $\sigma_{S}^{F}$ and $\sigma_{R}^{F}$ the corresponding standard deviations. In theory, $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ are equal to the ”true” transition probability (see Sect. 33332.2), but since $N$ and $M$ are finite, they may be biased. It is thus important to ensure that they lie within each other’s confidence interval. We measure the consistency of both distributions by comparing the distance between $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ with the overlap of the $95\%$ confidence intervals attached to both:
@@
where $t$ is the two-sided $95\%$ value from a Student’s distribution. A value of $\mathcal{C}<1$ ensures that both probability estimates are within the confidence interval of each other and thus that no method is biased (Rolland, 2022).
Figure 2: (a) Reproduction of the main result of Castellana et al. (2019). It is a diagram of transition probabilities depending on the parameters $(\overline{E_{a}},f_{\sigma})$ of the model. This diagram was computed with TAMS-$R$.
@@
The consistency indicator $\mathcal{C}$ is shown in Fig. 2b. $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ are consistent for most values of $(\overline{E_{a}},f_{\sigma})$ since $\mathcal{C}\geq 1$ only in $6.1\%$ of the cases. Fig. 2b can be divided into four regions. In the upper right corner, corresponding to large values of $\overline{E_{a}}$ and $f_{\sigma}$, the transition probabilities are close to $1$ and $\mathcal{C}$ is consistently below $10^{-2}$, which means that $\alpha_{S}^{F}$, $\alpha_{R}^{F}$ and their confidence intervals are almost identical. Then, for lower values of $\overline{E_{a}}$ and $f_{\sigma}$, $\alpha_{R}^{F}$ and $\alpha_{R}^{S}$ are still consistent but the blue show that the confidence intervals do not overlap perfectly anymore. However, $|\alpha_{R}^{F}-\alpha_{R}^{S}|$ remains approximately $10$ times smaller than the overlap of their confidence intervals, so these discrepancies are not significant. Then, the left part of the diagram, whatever the value of $f_{\sigma}$, is dominated by a region where $|\alpha_{S}^{F}-\alpha_{R}^{F}|$ is of the same order of magnitude as the overlap of their confidence interval, although most probability estimates remain consistent. This shows that as $\alpha_{R}^{F}$ decreases, the task of the RC becomes increasingly difficult and the transfer learning process propagates and amplifies biases. Finally, for $f_{\sigma}\leq 0.13$ and the smallest values of $\overline{E_{a}}$ such that $\alpha_{R}^{F}\geq 10^{-9}$, there is a region where inconsistencies dominate. In this regime of extremely low transition probabilities, TAMS-$R$ converges very slowly due to the biases propagated by the very long transfer learning, leading to largely underestimated values of $\alpha_{R}^{F}$. It is striking to note that only $4.5\%$ of the probability estimates $\alpha_{S}^{F}\geq 10^{-6}$ are inconsistent, but this number increases to $34.5\%$ for $\alpha_{S}^{F}<10^{-6}$. However, this also means that the TAMS-$R$ method is very reliable for transition probabilities down to $10^{-6}$, which is lower than often realistically needed.
@@ 3.2 Individual variances
$\sigma^{2}_{\mathrm{id}}$ (see equation (10)) is the optimal variance that a given TAMS result must have if $\Phi$ is the committor. If TAMS-$S$ or TAMS-$R$ has provided a certain transition probability, $(\sigma^{2})_{\mathrm{id},\{R,S\}}^{F}$ gives the variance associated with this exact output if it had been obtained with the committor function. So to compare the quality of TAMS-$S$ and TAMS-$R$, we compare the relative difference of their variance to their corresponding ideal variance:
@@
$V_{\mathrm{diff}}$ is thus a measure of the proximity to the committor: the smaller $V_{\mathrm{diff}}$, the closer $\Phi_{R}^{F}$ or $\Phi_{S}^{S}$ to the committor.
Since the results of TAMS-$R$ are averaged over $30$ independent RCs, to test the variance of this method accurately, we freeze the training of the $30$ RCs for some parameters $(\overline{E_{a}},f_{\sigma})$ and run TAMS-$R$ $30$ times for each of these fixed RCs. We then obtain $30$ variances around the mean probability estimates that are directly comparable to $(\sigma^{2})_{S}^{F}$. We do this for three values of noise amplitude: $f_{\sigma}=0.5,0.25$ and $0.05$. For each of these values of $f_{\sigma}$, we select equally spaced values of $\overline{E_{a}}$ ($1$ out of $10$ in the original $100$ used to compute 2a to explore the entire range of transition probabilities, from $10^{-9}$ to $1$). $V_{\mathrm{diff}}$ is shown in Fig. 3 for the three values of $f_{\sigma}$. On every panel, the blue line corresponds to $V_{\mathrm{diff},S}$ plotted against $\alpha_{S}^{F}$. Each thin black line represents $30$ runs of TAMS-$R_{i}$, where $i\in[1,30]$ represents each independent RC. Thus, each thin black line shows $V_{\mathrm{diff},R_{i}}$ plotted against $\alpha_{R_{i}}^{F}$. The thick black line is the average of the $30$ thin black lines.
@@
For all values of $f_{\sigma}$, there is a great variability between the RCs. However, for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$, despite this variability, all RCs are overall consistent with the behavior of $\Phi^{F}_{S}$. This can also be seen with the average of all RCs, which is very consistent with the curve corresponding to $\Phi^{F}_{S}$, both in terms of the mean transition probability (to be read in abscissa) and of the mean distance to the ideal score (to be read in ordinate). However, $\Phi^{F}_{S}$ systematically shows better agreement with the committor function because $V_{\mathrm{diff},S}$ is generally lower.
An interesting case is that of $f_{\sigma}=0.05$ (Fig. 3c). There, all RCs show a significant deviation from the behavior of $\Phi^{F}_{S}$. They also have much greater variability between them than for larger noise values, both in terms of distance from $(\sigma^{2})_{\mathrm{id},R_{i}}^{2}$ and of mean transition probability. The latter also deviates from what is obtained with $\Phi^{F}_{S}$, although both quantities remain consistent on average. This deviation illustrates the limits of transfer learning: the smaller $f_{\sigma}$, the longer the RC has been trained by transferring information learned during easier TAMS runs. When this process takes too long, the RC may be saturated with information and has difficulty adapting to new and more complex situations. Moreover, since all RCs are trained independently, the longer the transfer learning process lasts, the more likely it is to transfer and amplify biases.
@@ 3.3 Mean first passage times and trajectories
Finally, it is also important to make sure that the transition paths are not biased. To do this, two measures are considered: the mean first-passage time (MFPT) and the trajectories obtained as output of TAMS. TAMS only simulates trajectories until they reach the target set, so the MFPT can be sampled by looking at the last time step of these trajectories.
@@
Ensembles of $1,000$ trajectories were saved for several values of $(\overline{E_{a}},f_{\sigma})$, for both TAMS-$S$ and TAMS-$R$, and the MFPTs were averaged over each ensemble. The resulting diagram corresponding to $\Phi^{F}_{S}$ is shown in Fig. 4a. Fig. 4b plots against $\alpha_{S}^{F}$ the relative difference of the MFPTs obtained with $\Phi_{R}^{F}$ and $\Phi^{F}_{S}$. We see in Fig. 4a that the MFPTs lie between $4.2$ model years for the largest value of $\overline{E_{a}}$ and $f_{\sigma}$ to $62$ model years for $\overline{E_{a}}=0.26\mathrm{\ Sv}$ and $f_{\sigma}=0.05$. The distribution of MFPT values shown in Fig. 4a is the one we could expect, with the shortest MFPT corresponding to the largest transition probabilities and long MFPTs associated with very low transition probabilities.
Fig. 4b shows that there is a very good agreement between the MFPT as predicted by $\Phi^{F}_{S}$ and $\Phi_{R}^{F}$. For $\alpha_{S}^{F}$ of the order of $1$, their relative error can reach $12\%$ (in absolute value). But in that case, the transition times are typically shorter than $10$ years, which is too short to consider such an error to be significant; it can simply be due to the stochastic generation of trajectories. For values of $\alpha_{S}^{F}$ down to $10^{-6}$, the relative error in the estimated transition times is lower than $5\%$, therefore both methods give very consistent results. There is only one outlier: there is a $25\%$ relative error on the longest transition time, which corresponds to an overestimation by $15$ years when using $\Phi_{R}^{F}$. In general, the use of $\Phi_{R}^{F}$ systematically overestimates the transition times (by $3$ to $10\%$) for $\alpha_{S}^{F}<10^{-6}$.
@@
The last important quality criterion of a TAMS experiment is the trajectories themselves. To do this, we compare the ensembles obtained for $(\overline{E_{a}},f_{\sigma})=(0.288\mathrm{\ Sv},0.05)$ when sampling $1,000$ trajectories from $30$ runs of TAMS-$S$ and $30$ runs of TAMS-$R$. For $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$ (with similar values of $\overline{E_{a}}$), the results are similar as discussed below (not shown). In this case, $\alpha_{S}^{F}=4\times 10^{-4}$. We have chosen these system parameters as a kind of ”worst-case scenario” because the TAMS problem is particularly difficult and the RC is more likely to be imprecise due to the repeated transfer learning processes. Moreover, due to the large variance in the MFPT of the trajectories in this ensemble, we only consider here the trajectories whose transition time is comprised between $90$ and $100$ model years. The average and the interval between the $5^{\mathrm{th}}$ and $95^{\mathrm{th}}$ percentiles of each variable are shown in Fig. 5. The average agreement between the trajectories generated using the score $\Phi_{S}^{F}$ and $\Phi_{R}^{F}$ is very good. For all variables in the model, their averages and the distribution of $90\%$ of their values (throughout the ensemble) closely overlap. These transitions take $90$ to $100$ model years to occur, but we can see on the trajectories that the system stays around its steady on-state for the first $60$ model years.
The physics of the F-transition can also be understood from these trajectories. The transition can only begin when the random freshwater forcing $E_{s}+E_{a}$ is able to overcome the salt advection through the northern subtropical gyre $r_{n}$. Once a random variation of the asymmetric freshwater forcing $E_{a}$ results in a net freshening of the Northern Atlantic (decrease in $S_{n}$), the strength of the downward transport $q_{N}\propto(S_{n}-S_{ts})D^{2}$ also decreases. The salt-advection feedback and the fixed volume of the northern box result in an accumulation of saline water in the tropical box. This accumulation also triggers an increase in $D$, thus an increase of $V_{ts}$ and a freshening of $S_{ts}$. However, this freshening is countered by the increasing effect of the southern subtropical gyre and of the wind-driven Ekman flow (due to the salinification of the tropics). As a result, $S_{ts}$ freshens slower than $S_{n}$, weakening even more $q_{N}$. The increase in $D$ and the effect of the northern tropical gyre are not sufficient to counteract the freshening of the Northern Atlantic and the downwelling stops when $S_{n}$ drops below a critical threshold. Figure 5 shows that during this process, only $S_{n}$, $S_{ts}$ and $D$ vary significantly, validating in hindsight our restriction to these variables for $\Phi_{R}^{F}$.

### v3

@@ 3 Results: F-transitions
In Castellana et al. (2019), the score function for TAMS is:
@@
where $\mathbf{x}\in\mathbb{R}^{d}$ is a state in the phase space and the exponent $ON$ refers to the steady on-state of the AMOC model. $\Phi^{F}_{S}(\mathbf{x})=1$ whenever $q_{N}(\mathbf{x})=0$. Given a certain parameter $\rho$, we define $A$ to be a neighborhood of $q^{ON}_{N}$ such that $q_{N}(\mathbf{x})\geq q_{S}(\mathbf{x})$: $\ A=\{\mathbf{x}\in\mathbb{R}^{n}\ |\ ||\mathbf{x}-\mathbf{x}^{ON}||\leq\rho\ \&\ q_{N}(\mathbf{x})\geq q_{S}(\mathbf{x})\}$. $\Phi^{F}_{S}(\mathbf{x})$ is set to 00 in this area.
The score function estimated by the RC is called $\Phi_{R}^{F}$. We call TAMS-$S$ the combination of TAMS with $\Phi^{F}_{S}$ and TAMS-$R$ the combination of TAMS with $\Phi_{R}^{F}$. We sample $100$ equally spaced values of $\overline{E_{a}}\in[0.06,0.35]\mathrm{\ Sv}$ and $f_{\sigma}\in[0.005,0.5]$. TAMS uses an ensemble of $N=1,000$ trajectories, $n_{c}=10$ of which are discarded at each iteration; $T_{\mathrm{max}}=100$ years. For each couple $(\overline{E_{a}},f_{\sigma})$, TAMS-$S$ and TAMS-$R$ are run both $M=30$ times. TAMS is stopped immediately after $\Phi^{F}_{\{R,S\}}=0$, before the circulation in the Southern Atlantic can be reversed, so that we are certain to sample only F-transitions with this method.
@@ 3.1 Consistency indicator $\mathcal{C}$
First, the transition probability estimates obtained with both methods must be consistent. We call $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ the mean probability estimates, respectively, obtained with TAMS-$S$ and TAMS-$R$. We call $\sigma_{S}^{F}$ and $\sigma_{R}^{F}$ the corresponding standard deviations. We measure the consistency of the distributions of $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ by comparing the difference of their mean values to the overlap of their $95\%$ confidence interval:
@@
where $t$ is the two-sided $95\%$ value from a Student’s distribution. $\mathcal{C}<1$ ensures that both probability estimates are within the confidence interval of each other so that no method is biased (Rolland, 2022).
Figure 2: (a) Estimated transition probabilities using TAMS-$R$ depending on the model parameters $(\overline{E_{a}},f_{\sigma})$. This reproduces the main result of Castellana et al. (2019), who used TAMS-$S$.
@@
Fig. 2b shows the consistency indicator $\mathcal{C}$. $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ are consistent for most values of $(\overline{E_{a}},f_{\sigma})$ since $\mathcal{C}\geq 1$ only in $6.1\%$ of the cases. Fig. 2b can be divided into four regions. In the upper right corner, corresponding to large values of $\overline{E_{a}}$ and $f_{\sigma}$, the transition probabilities are close to $1$ and $\mathcal{C}$ is consistently below $10^{-2}$, so $\alpha_{S}^{F}$, $\alpha_{R}^{F}$ and their confidence intervals are almost identical. For lower values of $\overline{E_{a}}$ and $f_{\sigma}$, $\alpha_{R}^{F}$ and $\alpha_{R}^{S}$ are still consistent but the lighter blue shading shows that the confidence intervals do not perfectly overlap anymore. However, $|\alpha_{R}^{F}-\alpha_{R}^{S}|$ remains approximately $10$ times smaller than the overlap of their confidence intervals, so these discrepancies are not significant. Then, whatever the value of $f_{\sigma}$, in the region containing the smallest values of $\overline{E_{a}}$ such that $\alpha_{R}^{F}>10^{-9}$, all values of $|\alpha_{S}^{F}-\alpha_{R}^{F}|$ are of the same order of magnitude as the overlap of their confidence interval, although most probability estimates remain consistent. This shows that as $\alpha_{R}^{F}$ decreases, the task of the RC becomes increasingly difficult and the transfer learning propagates and amplifies biases. Finally, for $f_{\sigma}\leq 0.13$ and the smallest values of $\overline{E_{a}}$ such that $\alpha_{R}^{F}\geq 10^{-9}$, there is a region where inconsistencies dominate. In this regime of extremely low transition probabilities, TAMS-$R$ converges very slowly due to the biases propagated by the very long transfer learning and largely underestimates $\alpha_{R}^{F}$. It is striking to note that only $4.5\%$ of the probability estimates $\alpha_{S}^{F}\geq 10^{-6}$ are inconsistent, but this number increases to $34.5\%$ for $\alpha_{S}^{F}<10^{-6}$. However, this also means that the TAMS-$R$ method is very reliable for transition probabilities down to $10^{-6}$, which is lower than often realistically needed. Collins et al. (2013) and Kriegler et al. (2009) estimate that the probability that this transition occurs within $100$ years has an upper bound of a few percent, which is many orders of magnitude above the lowest probability that our method can estimate. Moreover, this probability increases with global warming (Kriegler et al., 2009; Armstrong McKay et al., 2022). Finally, studies (Hall and Ronald, 2001; Li and Born, 2019; Vettoretti et al., 2022) have shown from paleoclimate simulations that such transition has a non-negligible probability, as it can occur spontaneously several times per $10^{4}$ years.
@@ 3.2 Individual variances
$(\sigma^{2})_{\mathrm{id},\{R,S\}}^{F}$ (see equation (9)) gives the optimal variance associated with the transition probability computed with TAMS-$S$ or TAMS-$R$ if it had been obtained using the committor function. So to compare the quality of TAMS-$S$ and TAMS-$R$, we compare the relative difference of their variance to their corresponding ideal variance:
@@
$V_{\mathrm{diff}}$ is a measure of the proximity to the committor: the smaller $V_{\mathrm{diff}}$, the closer $\Phi_{R}^{F}$ or $\Phi_{S}^{S}$ to the committor.
In Fig. 2a, TAMS-$R$ was run $30$ times using $30$ independent RCs acting as $30$ different score functions. The variance over these runs cannot be directly compared to the variance over the $30$ runs of TAMS-$S$, which uses a fixed score function. When running TAMS-$R$, the RCs are updated at each TAMS iteration, adapting them to each couple of parameters $(\overline{E_{a}},f_{\sigma})$. These trained RCs can then be “frozen” (their parameters are fixed and not retrained anymore) and be used as fixed score function for TAMS-$R$, so that the variance on the TAMS results when using RCs can be compared to the variance of TAMS-$S$. TAMS-$R$ is thus run $30$ times again using each of the $30$ independent RCs trained previously ($900$ runs of TAMS in total). We do this for three values of noise amplitude: $f_{\sigma}=0.5,0.25$ and $0.05$. For each of these values of $f_{\sigma}$, results are saved for one value of $\overline{E_{a}}$ out of $10$ across the whole range $[0.06,0.35]\mathrm{\ Sv}$ (the computation is stopped when $\alpha_{R}^{F}\leq 10^{-9}$). $V_{\mathrm{diff}}$ is shown in Fig. 3 for the three values of $f_{\sigma}$. On every panel, the blue line corresponds to $V_{\mathrm{diff},S}$ plotted against $\alpha_{S}^{F}$. Each thin black line represents $30$ runs of TAMS-$R_{i}$, where $i\in[1,30]$ represents each independent RC. Thus, each thin black line shows $V_{\mathrm{diff},R_{i}}$ plotted against $\alpha_{R_{i}}^{F}$. The thick black line is the average of the $30$ thin black lines.
@@
For all values of $f_{\sigma}$, there is a great variability between the RCs. However, for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$, all RCs are overall consistent with $\Phi^{F}_{S}$. The average of all RCs is also consistent with the curve corresponding to $\Phi^{F}_{S}$, both in terms of the mean transition probability (to be read in abscissa) and of the mean distance to the ideal score (to be read in ordinate). However, $\Phi^{F}_{S}$ shows better agreement with the committor function as $V_{\mathrm{diff},S}$ is generally lower.
For $f_{\sigma}=0.05$ (Fig. 3c), all RCs show a significant deviation from $\Phi^{F}_{S}$ and have a much greater variability between them in terms of distance to $(\sigma^{2})_{\mathrm{id},R_{i}}^{F}$ than for larger $f_{\sigma}$. The lowest values of $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ also deviate (the left-end of the curves are not at the same point on the horizontal axis), although they remain consistent on average (thick black and blue lines) at about $10^{-8}$. This deviation illustrates the limits of transfer learning: the smaller $f_{\sigma}$, the longer the RC was trained by transferring information learned during easier TAMS runs. When this process takes too long, the RC may be saturated with information and has difficulty adapting to new and more complex situations. Moreover, since all RCs are independent, the longer the transfer learning process, the more likely it is to amplify biases.
@@ 3.3 Mean first passage times and trajectories
Finally, the transition paths must not be biased. Two measures are considered: the mean first-passage time (MFPT) and the trajectories obtained as output of TAMS. TAMS only simulates trajectories until they reach the target set, so the MFPT can be estimated from the last time step of these trajectories.
@@
Ensembles of $1,000$ trajectories were saved for several values of $(\overline{E_{a}},f_{\sigma})$, for both TAMS-$S$ and TAMS-$R$, and the MFPTs were averaged over each ensemble. The resulting diagram corresponding to $\Phi^{F}_{S}$ is shown in Fig. 4a. Fig. 4b plots against $\alpha_{S}^{F}$ the relative difference of the MFPTs obtained with $\Phi_{R}^{F}$ and $\Phi^{F}_{S}$. Fig. 4a shows that the MFPTs lie between $4.2$ model years for the largest value of $\overline{E_{a}}$ and $f_{\sigma}$ to $62$ model years for $\overline{E_{a}}=0.26\mathrm{\ Sv}$ and $f_{\sigma}=0.05$. This distribution of MFPTs is as expected, with the shortest MFPTs corresponding to the largest transition probabilities and long MFPTs associated with low transition probabilities.
Fig. 4b shows that the MFPTs predicted by $\Phi^{F}_{S}$ and $\Phi_{R}^{F}$ are in agreement. For $\alpha_{S}^{F}$ of the order of $1$, their relative error can reach $12\%$ (in absolute value). But in that case, transition times are typically shorter than $10$ years, so such an error can simply be due to the stochastic generation of trajectories. For values of $\alpha_{S}^{F}$ down to $10^{-6}$, the relative error in the estimated transition times is lower than $5\%$, therefore both methods are very consistent. The only outlier is the $25\%$ relative error on the longest transition time, which corresponds to an overestimation by $15$ years when using $\Phi_{R}^{F}$. In general, the use of $\Phi_{R}^{F}$ systematically overestimates the transition times (by $3$ to $10\%$) for $\alpha_{S}^{F}<10^{-6}$.
@@
Then, we compare the trajectory ensembles obtained for $(\overline{E_{a}},f_{\sigma})=(0.288\mathrm{\ Sv},0.05)$ when sampling $1,000$ trajectories from $30$ runs of TAMS-$S$ and $30$ runs of TAMS-$R$. For $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$ (with similar values of $\overline{E_{a}}$), the results are similar as discussed below (not shown). In this case, $\alpha_{S}^{F}=4\times 10^{-4}$. We have chosen these system parameters as a ”worst-case scenario” because the TAMS problem is particularly difficult and the RC is more likely to be imprecise due to the repeated transfer learning processes. Moreover, due to the large variance in the MFPTs of the trajectories in this ensemble, we only consider here the trajectories whose transition time lies between $90$ and $100$ model years. This corresponds to $124$ trajectories for TAMS-$S$ and $145$ for TAMS-$R$ because, in TAMS, trajectories are left free to collapse at any time during the $100$ simulated years. We restricted ourselves to this decade to show typical collapses from trajectories within a full century. The average and the interval between the $5^{\mathrm{th}}$ and $95^{\mathrm{th}}$ percentiles of each variable are shown in Fig. 5. The average agreement between the trajectories generated using the score $\Phi_{S}^{F}$ and $\Phi_{R}^{F}$ is very good. For all variables in the model, their averages and the distribution of $90\%$ of their values (throughout the ensemble) closely overlap. These transitions take $90$ to $100$ model years to occur, but we can see on the trajectories that the system stays around its steady on-state for the first $60$ model years.